# Does Your Multi-modal LLM Truly See the Diagrams in Visual Math Problems?

### v1

@@ F.2 Response of Different MLLMs
Figure 23: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 24: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 25: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 26: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 27: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 28: Response Comparison of GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
@@ F.3 Response of Different Problem Versions
Figure 29: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 30: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 31: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 32: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 33: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 34: Response of Different Problem Versions by GPT-4V openai2023gpt4v . By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.

### v2

@@ F.2 Response of Different MLLMs
Figure 23: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 24: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 25: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 26: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 27: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
Figure 28: Response Comparison of GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. We adopt the Text-lite version of the problem, and highlight the key-step extraction and scoring by the CoT evaluation strategy.
@@ F.3 Response of Different Problem Versions
Figure 29: Response of Different Problem Versions by GPT-4V [47]. By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 30: Response of Different Problem Versions by GPT-4V [47]. By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 31: Response of Different Problem Versions by GPT-4V [47]. By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 32: Response of Different Problem Versions by GPT-4V [47]. By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 33: Response of Different Problem Versions by GPT-4V [47]. By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.
Figure 34: Response of Different Problem Versions by GPT-4V [47]. By comparing the output from three versions, we observe increasing reasoning errors when more textual content is visually depicted in diagrams. We only showcase the Text-dominant version of the problem for brevity.